Consider the following reaction:
2NiO(s)→2Ni(s)+O2(g)

Vapor Pressure of Water versus Temperature
Temperature
(∘C)
Pressure
(mmHg)
Temperature
(∘C)
Pressure
(mmHg)
0 4.58 55 118.2
5 6.54 60 149.6
10 9.21 65 187.5
15 12.79 70 233.7
20 17.55 75 289.1
25 23.78 80 355.1
30 31.86 85 433.6
35 42.23 90 525.8
40 55.40 95 633.9
45 71.97 100 760.0
50 92.6

Part A

If O2 is collected over water at 40.0 ∘C and a total pressure of 742 mmHg , what volume of gas will be collected for the complete reaction of 25.09 g of NiO?

V

V

=
L
Answer #1

Balanced equation:
2 NiO(s) ====> 2 Ni(s) + O2(g)

Mass of NiO = 25.09 gm

Molar mass of NiO = 74.69 g/mol

Moles of NiO = 25.09 g / 74.69 g /mol = 0.3359 Moles

Moles of O2 produced =  0.1679 Moles

Using the mole and following formula we can calculate the volume of oxygen.

PV= nRT

P = Pressure in atm                       V= Volume in Liter

n = no of moles               R = 0.08206 L atm K-1 Mol-1

T = Temperature in Kelvin

P = 742 - 55.4 = 686.6 mmHg = 0.90342 atm V = ? T = 273 + 40 = 313 K n = 0.1679 Moles

V = nRT / P = 0.1679 Moles x 0.08206 L atm K-1 Mol-1 x 313 K / 0.90342 atm

V = 4.7734 Liter

Hence 4.7734 Liter of O2 gas will be produced.